# Learning to Think: Information-Theoretic Reinforcement Fine-Tuning for LLMs

### v1

@@ 3.3 Motivation Analysis
First, recall the limitations of existing methods that analyzed above: they rely on outcome rewards and extend the reasoning chain to fully explore the solution space, securing reasoning performance; however, Subsection 3.2 demonstrates that they may not efficiently leverage test-time compute and squanders resources. To balance reasoning effectiveness and efficiency, we aim for an algorithm that yields positive gains at every reasoning step, enabling the model to achieve comparable reasoning performance within a constrained yet sufficient token budget. This requires augmenting the learning objective with dense, step-wise process rewards that immediately quantify each reasoning step’s contribution to overall performance. By maximizing the cumulative reward across all reasoning steps, we encourage the model to generate tokens that most benefit the answer correctness. Therefore, the algorithm requires an effective dense process reward to ensure both effectiveness and efficiency: by maximizing total reward across all reasoning steps, we guarantee the answer accuracy; by maximizing each token’s contribution and preventing useless tokens, we ensure efficiency. Notably, some concurrent approaches attempt to reduce reasoning depth using task-specific process rewards zhang2024rest ; zhou2025sweetrltrainingmultiturnllm and length penalties ye2025emergence ; luo2025o1 ; however, they require manually designing high-quality process labels and task-specific evaluators, which is an expensive endeavor that may not generalize across tasks. Therefore, we must also address the second challenge of algorithmic generality.
@@ 4 Learning To Think
Building on the above insights, we propose Learning to Think (L2T), an information-theoretic reinforcement fine-tuning framework for LLMs (with pseudo-code in Appendix B). The key is proposing a universal dense process reward for LLM optimization to adaptively allocate reasoning depth across tasks and prevent token waste. It recasts LLM optimization as an episodic learning problem to support episode-wise process reward calculation and optimizes LLM to maximize the contribution of tokens generated in each episode, ensuring accuracy and efficiency. Specifically, in L2T, each query–response pair is segmented into successive reasoning episodes (Subsection 4.1), within which the model performs an adaptation update to increase the likelihood of a correct answer. To ensure the effectiveness of each adaptation and curb excessive reasoning, we then propose an information-theoretic dense process reward that immediately quantifies the progress of each episode with universal information gain to support policy optimization (Subsection 4.2). It can be efficiently estimated leveraging PAC-Bayes bounds and the Fisher information matrix with theoretical guarantees (Theorem 4.2). Finally, we optimize the LLMs by maximizing cumulative reward via RL (Subsection 4.3), ensuring both high reasoning accuracy and efficiency across different tasks.
@@ 4.1 Problem Reformulation
In practice, reasoning questions vary widely, e.g., from simple arithmetic to complex mathematical proofs. Existing RL-based fine-tuning paradigm wiering2012reinforcement ; szepesvari2022algorithms mainly treats the reasoning process of each question as a single episode and updates the policy via the final outcome reward. Under this setting, it is difficult to implement dense process rewards. To address this issue, we reformulate the problem mentioned in Subsection 3.1 as an episodic RL problem. We treat each question as a task sampled from a broader distribution and decompose the reasoning process into successive episodes. This decomposition allows us to assign process rewards at every intermediate episode, with each episode’s reward reflecting its contribution to the answer correctness. Based on this dense reward, we can optimize the policy incrementally to maximize the performance gain of each episode.
@@
where $B_{token}$ is the fixed test-time compute budget. Through Eq.2, the learned policy allocates its limited token budget ($B_{token}$) where it yields the greatest incremental benefit, expanding promising lines with higher process reward $r^{\rm prg}$ and maximize overall task success $r^{\rm out}$. In practice, however, the crux is the design of the newly introduced dense process reward $r^{\rm prg}$, which need to satisfy following desiderata: (i) Relevance: faithfully measure the progress that episode $k$ contributes toward a correct solution; (ii) Efficiency: be cheap to compute; (iii) Generality: apply uniformly across tasks without bespoke engineering for each new domain. Recently proposed process-reward models zhang2024entropyregularizedprocessrewardmodel ; qu2025optimizing ; zhang2024rest ; zhou2025sweetrltrainingmultiturnllm predominantly focus on (i); however, they remain task-specific and depend on high-quality annotations, failing to satisfy (ii) and (iii). Therefore, designing such a dense, per-episode reward remains an open challenge, also the key to unlocking truly efficient reasoning in LLMs.
@@ 4.2 Learning Information-Theoretic Dense Process Reward
To address the above challenge, we propose a novel information-theoretic dense process reward. It is inspired by the information theory ash2012information ; khinchin2013mathematical and consists of two constraints for reasoning effectiveness and efficiency: (i) a fitting information gain, which encourages the model to acquire key information about correctness in each episode; and (ii) a parameter compression penalty, which penalizes redundant information absorbed at each episode to maintain efficiency. This reward is agnostic to input format or task domain, which can be applied in various scenarios. In this subsection, we begin by explaining the meaning of the above two constraints with the proposed reward. Then, we provide the formal definition of this reward (Definition 4.1). Next, we explain how to efficiently compute it in practice: to handle the large parameter scale of LLMs, we develop an efficient approximation of this reward using PAC-Bayes theory and Fisher information matrix (Theorem 4.2). Finally, we illustrate why the proposed reward is effective, i.e., satisfying the three criteria mentioned in Subsection 4.1.
Firstly, we explain what the two components of the proposed reward are. Specifically, the fitting information gain measures the reduction in uncertainty about $Y$ given $X$ provided by parameters $\theta$ after each episode; formally, it is defined as the conditional mutual information $I(\theta;Y|X)=H(Y|X)-H(Y|X,\theta)$, where $H(Y|X)=-\sum_{y}p(y|X)\log p(y|X)$ denotes the uncertainty of $Y$ given $X$ alone and $H(Y|X,\theta)$ the residual uncertainty after observing $\theta$. The fitting gain for episode $k$, in which parameters update from $\theta_{k-1}$ to $\theta_{k}$, is $\Delta I_{k}=I(\theta_{k};Y|X)-I(\theta_{k-1};Y|X)$. Since direct computation is costly in LLMs, we approximate $\Delta I_{k}$ by the increase in the predicted correctness probability of the models, i.e., $\Delta I_{k}\approx J_{r}(\pi_{\theta}(\cdot|s_{k},z_{k}))-J_{r}(\pi_{\theta}(\cdot|s_{k}))$, which aligns with the direction of $\Delta I_{k}$ (Appendix C.1). It entails just two forward passes of $\pi_{\theta}$ without the need of gradient updates to estimate $\Delta I_{k}$, reducing computational overhead. The parameter compression penalty constrains redundant information captured from each episode, preventing excessive updates. It is defined as the mutual information between $\theta$ and the context $s_{k}$, denoted as $I(\theta;s_{k})=\mathbb{E}_{S}[\mathrm{KL}(p(\theta|s_{k})\,\|\,p(\theta))]$. It quantifies the task-specific idiosyncrasies stored in $\theta$ gray2011entropy , where larger mutual information implies greater overfitting risk and unnecessary computational overhead. Thus, these terms align with our objective of evaluating reasoning effectiveness and efficiency in dense process reward design.
@@
where $J_{r}(\cdot)$ denotes the correctness prediction probability and $\beta>0$ is a hyperparameter.
@@
Obtaining Definition 4.1, we illustrate how to calculate it in practice. In the LLM setting, fitting information gain measures the contribution of each optimization episode to reasoning capability by tracking the change in the predicted correctness probability, i.e., the output distributions of $\pi_{\theta}$. In contrast, computing the parameter-compression penalty is more involved: it requires quantifying the mutual information increment between the parameters $\theta$ and the historical context $s_{k}$, where direct estimation in the large parameter space of LLMs is intractable. To address this, we introduce an efficient approximation that uses the low-rank parameter proxy $\tilde{\theta}$ with singular value decomposition (SVD) brand2006fast and the Fisher information matrix frieden2004science ; reza1994introduction to estimate the penalty term. We get:
Theorem 4.2
@@
Theorem 4.2 presents a method to estimate the intractable compression penalty (with proof in Appendix A). It simplifies the parameter space using SVD and assumes $\tilde{\theta}$ follows a Gaussian distribution based on kwak2017central , a common and mild assumption. By computing the second derivative of the log-likelihood of $\theta$, we obtain the Fisher information matrix, which captures the effect of $\tilde{\theta}_{k}$ updates on the output and approximates the covariance calculation (Lemma A.1). The mutual information increment is then approximated using the second-order term of the Taylor expansion. This method significantly reduces the computational complexity (Theorem C.2) with limited approximation error (Theorem C.3) to support the computation of our proposed reward Definition 4.1 in practice.
@@ 4.3 Optimizing LLM with Reinforcement Fine-Tuning
where $\alpha$ is the importance weight (set to 1 for simplicity), and the $r_{k}^{\rm prg}$ is calculated through Theorem 4.2. Through Eq.5, L2T optimizes the policy $\pi_{\theta}$ to achieve our goal of boosting effectiveness and efficiency in two parts: (i) The first part ensures the correctness of the final answer through the outcome reward $r(x,z_{0:K})$. (ii) The second part introduces the dense process reward $r_{k}^{\rm prg}$, leveraging information-theoretic internal signals to assess the progress of each episode update. It encourages the model to maximize correctness at each step while avoiding redundant information accumulation. Thus, this optimization enables the model to efficiently utilize the limited token budget, progressively improving reasoning results, and ultimately achieving high-accuracy reasoning outputs.
@@ 5.1 Experimental Settings
We evaluate on multiple reasoning benchmarks, including AIME24–25, AMC, MATH500 hendrycks2021measuring , MinervaMATH lewkowycz2022solving , and Omni-MATH gao2024omnimathuniversalolympiadlevel (see Appendices D and F for more benchmarks, e.g., code generation). We use DeepScaleR-1.5B-Preview and DeepSeek-R1-Distill-Qwen-1.5B as base models, which already generate reasoning traces marked with ‘<think>’. We compare L2T against (i) outcome reward-based RL methods (e.g., GRPO shao2024deepseekmathpushinglimitsmathematical for deepseek-model family, more in appendices) and (ii) test-time-compute–focused methods, e.g., length penalty arora2025training and process-reward approaches such as ReST-MCTS zhang2024rest and MRT qu2025optimizing . Since DeepScaleR-1.5B-Preview has already undergone one round of fine-tuning on 40k math question–answer pairs, we fine-tune it on the 919 AIME questions (from 1989 to 2023); for DeepSeek-R1-Distill-Qwen-1.5B, we fine-tune on a random sample of 4,000 question–answer pairs from NuminaMath li2024numinamath . Both fine-tuning and evaluation use a maximum token budget of 16,384. For optimization, we set the learning rate to $1e^{-6}$, weight decay to 0.01, and batch size to 256. To approximate the parameter-compression penalty, we employ a one-layer MLP with a Fisher information–matrix damping coefficient of $1\times 10^{-5}$. The hyperparameters $\alpha$ (Eq.5) and $\beta$ (Eq.3) are set to 0.8 and 0.6, respectively. All experiments are run on the A100 GPU clusters. More details of implementation and hyperparameters are provided in Appendix E.
@@
+outcome-reward RL (GRPO) 44.5 (+1.7) 39.3 (+2.6) 81.5 (–1.5) 84.9 (-0.3) 24.7 (+0.1) 55.0 (+0.5)
+length penalty 40.3 (–2.5) 30.3 (–6.4) 77.3 (–5.7) 83.2 (-2.0) 23.0 (-1.6) 50.8 (–3.7)
@@ 5.2 Effectiveness and Efficiency Analysis
We compare L2T with all the baselines across all the benchmarks and base models, recording both pass@1 accuracy and the required token budget. To reduce variance from limited samples, we use 20 outputs per question. Table 1 and Figure 2 show that L2T attains SOTA performance, achieving the highest reasoning accuracy with the smallest token budget. For example, compared to outcome-reward-based method, L2T delivers over a 10% gain in pass@1 and roughly doubles token efficiency; compared to methods focused on test-time compute, it achieves more than a 5% accuracy improvement while reducing the token budget by 20%. Moreover, L2T consistently outperforms baselines on multiple datasets with distributions different from the training data, further demonstrating its effectiveness across diverse reasoning tasks. These results confirm the superiority of our approach, which achieves better reasoning with higher efficiency.
@@ 5.3 Ablation study
The effect of different components. We evaluate three alternative configurations: (i) replacing the fitting information gain with a task-specific reward model; (ii) removing the parameter-compression penalty; and (iii) substituting the low-rank approximation with random sampling of 30% layers. Notably, the overall contribution of our reward has already been demonstrated in Subsection 5.2. From Figure 5 and Appendix F.3, we observe that both the fitting information gain and the parameter-compression penalty are critical for LLM reasoning; although random sampling is faster than low-rank approximation, it introduces additional error. These findings underscore the advantages of our design.
Parameter sensitivity. We determine the hyperparameters of L2T by evaluating reasoning performance across benchmarks. Both $\alpha$ and $\beta$ are swept over $[0.3,0.9]$. We first use grid search to screen the parameters with a difference of 0.05, then refine it with 0.01, recording the average outcome. As shown in Figure 5, the optimal setting is $\alpha=0.8$ and $\beta=0.6$, also our choices.

### v2

@@ 3.3 Motivation Analysis
As illustrated in Section 1 and Subsection 3.2, we discuss and demonstrate that, although outcome rewards can boost reasoning effectiveness by increasing the token budget, it sacrifices efficiency, i.e., many reasoning steps are invalid for the answer correctness; much longer CoTs are used than the correct answer really needed. To balance reasoning effectiveness and efficiency, we aim for an algorithm that yields positive gains at every reasoning step, enabling the model to achieve comparable reasoning performance within a constrained yet sufficient token budget. This requires augmenting the learning objective with dense, step-wise process rewards that immediately quantify each reasoning step’s contribution to overall performance. By maximizing the cumulative reward across all reasoning steps, we encourage the model to generate tokens that most benefit the answer correctness. Therefore, the algorithm requires an effective dense process reward to ensure both effectiveness and efficiency: by maximizing total reward across all reasoning steps, we guarantee the reasoning effectiiveness; by maximizing each token’s contribution and preventing useless tokens, we ensure efficiency. Notably, some concurrent approaches attempt to reduce reasoning depth using task-specific process rewards [62, 67] and length penalties [57, 33]; however, they require manually designing high-quality process labels and task-specific evaluators, which is an expensive endeavor that may not generalize across tasks. Therefore, we must also address the second challenge of algorithmic generality.
@@ 4 Learning To Think
Based on the above analyses, we propose Learning to Think (L2T), an information-theoretic reinforcement fine-tuning framework for LLMs (with pseudo-code in Appendix B). The key is proposing a universal dense process reward for LLM optimization to adaptively allocate reasoning depth across tasks and prevent token waste. It recasts LLM optimization as an episodic learning problem to support episode-wise process reward calculation and optimizes LLM to maximize the contribution of tokens generated in each episode, ensuring reasoning effectiveness and efficiency. Specifically, in L2T, each query-response pair is segmented into successive reasoning episodes (Subsection 4.1), within which the model performs an adaptation update to increase the likelihood of a correct answer. To ensure the effectiveness of each adaptation and curb excessive reasoning, we then propose an information-theoretic dense process reward that immediately quantifies the progress of each episode with universal information gain to support policy optimization (Subsection 4.2). It can be efficiently estimated leveraging PAC-Bayes bounds and the Fisher information matrix with theoretical guarantees (Theorem 4.2). Finally, we optimize the LLMs by maximizing cumulative reward via RL (Subsection 4.3), ensuring both high reasoning effectiveness and efficiency across different tasks.
@@ 4.1 Problem Reformulation
In practice, reasoning questions vary widely, e.g., from simple arithmetic to complex mathematical proofs. Existing RL-based fine-tuning paradigm [51, 45] mainly treats the reasoning process of each question as a single episode and updates the policy via the final outcome reward. Under this setting, it is difficult to implement dense process rewards. To address this issue, we reformulate the problem mentioned in Subsection 3.1 as an episodic RL problem. We treat each question as a task sampled from a broader distribution and decompose the reasoning process into successive episodes. This decomposition allows us to assign process rewards at every intermediate episode, with each episode’s reward reflecting its contribution to the answer correctness. Based on this dense reward, we can optimize the policy incrementally to maximize the performance gain of each episode.
@@
where $B_{token}$ is the fixed test-time compute budget. Through Eq.2, the learned policy allocates its limited token budget ($B_{token}$) where it yields the greatest incremental benefit, expanding promising lines with higher process reward $r^{\rm prg}$ and maximize overall task success $r^{\rm out}$. In practice, however, the crux is the design of the newly introduced dense process reward $r^{\rm prg}$, which need to satisfy following desiderata: (i) Relevance: faithfully measure the progress that episode $k$ contributes toward a correct solution; (ii) Efficiency: be cheap to compute; (iii) Generality: apply uniformly across tasks without bespoke engineering for each new domain. Recently proposed process-reward models [63, 36, 62, 67] predominantly focus on (i); however, they remain task-specific and depend on high-quality annotations, failing to satisfy (ii) and (iii). Therefore, designing such a dense, per-episode reward remains an open challenge, also the key to unlocking truly efficient reasoning in LLMs.
@@ 4.2 Learning Information-Theoretic Dense Process Reward
To address the above challenge, we propose a novel information-theoretic dense process reward. It is inspired by the information theory [3, 25] and consists of two constraints for reasoning effectiveness and efficiency: (i) a fitting information gain, which encourages the model to acquire key information about correctness in each episode; and (ii) a parameter compression penalty, which penalizes redundant information absorbed at each episode to maintain efficiency. This reward is agnostic to input format or task domain, which can be applied in various scenarios. In this subsection, we begin by explaining the meaning of the above two constraints with the proposed reward. Then, we provide the formal definition of this reward (Definition 4.1). Next, we explain how to efficiently compute it in practice: to handle the large parameter scale of LLMs, we develop an efficient approximation of this reward using PAC-Bayes theory and Fisher information matrix (Theorem 4.2). Finally, we illustrate why the proposed reward is effective, i.e., satisfying the three criteria mentioned in Subsection 4.1.
Firstly, we explain what the two components of the proposed reward are. Specifically, the fitting information gain measures the reduction in uncertainty about $Y$ given $X$ provided by parameters $\theta$ after each episode; formally, it is defined as the conditional mutual information $I(\theta;Y|X)=H(Y|X)-H(Y|X,\theta)$, where $H(Y|X)=-\sum_{y}p(y|X)\log p(y|X)$ denotes the uncertainty of $Y$ given $X$ alone and $H(Y|X,\theta)$ the residual uncertainty after observing $\theta$. The fitting gain for episode $k$, in which parameters update from $\theta_{k-1}$ to $\theta_{k}$, is $\Delta I_{k}=I(\theta_{k};Y|X)-I(\theta_{k-1};Y|X)$. Since direct computation is costly in LLMs, we approximate $\Delta I_{k}$ by the increase in the predicted correctness probability of the models, i.e., $\Delta I_{k}\approx J_{r}(\pi_{\theta}(\cdot|s_{k},z_{k}))-J_{r}(\pi_{\theta}(\cdot|s_{k}))$, which aligns with the direction of $\Delta I_{k}$ (Appendix C.1). It entails just two forward passes of $\pi_{\theta}$ without the need of gradient updates to estimate $\Delta I_{k}$, reducing computational overhead. The parameter compression penalty constrains redundant information captured from each episode, preventing excessive updates. It is defined as the mutual information between $\theta$ and the context $s_{k}$, denoted as $I(\theta;s_{k})=\mathbb{E}_{S}[\mathrm{KL}(p(\theta|s_{k})\,\|\,p(\theta))]$. It quantifies the task-specific idiosyncrasies stored in $\theta$ [14], where larger mutual information implies greater overfitting risk and unnecessary computational overhead. Thus, these terms align with our objective of evaluating reasoning effectiveness and efficiency in dense process reward design.
@@
where $J_{r}(\cdot)$ denotes the correctness probability and $\beta>0$ is a hyperparameter.
@@
Obtaining Definition 4.1, we illustrate how to calculate it in practice. In the LLM setting, fitting information gain measures the contribution of each optimization episode to reasoning capability by tracking the change in the predicted correctness probability, i.e., the output distributions of $\pi_{\theta}$. In contrast, computing the parameter-compression penalty is more involved: it requires quantifying the mutual information increment between the parameters $\theta$ and the historical context $s_{k}$, where direct estimation in the large parameter space of LLMs is intractable. To address this, we introduce an efficient approximation that uses the low-rank parameter proxy $\tilde{\theta}$ with singular value decomposition (SVD) [6] and the Fisher information matrix [12, 37] to estimate the penalty term. We get:
Theorem 4.2.
@@
Theorem 4.2 presents a method to estimate the intractable compression penalty (with proof in Appendix A). It simplifies the parameter space using SVD and assumes $\tilde{\theta}$ follows a Gaussian distribution based on [26], a common and mild assumption. By computing the second derivative of the log-likelihood of $\theta$, we obtain the Fisher information matrix, which captures the effect of $\tilde{\theta}_{k}$ updates on the output and approximates the covariance calculation (Lemma A.1). The mutual information increment is then approximated using the second-order term of the Taylor expansion. This method significantly reduces the computational complexity (Theorem C.2) with limited approximation error (Theorem C.3) to support the computation of our proposed reward Definition 4.1 in practice.
@@ 4.3 Optimizing LLM with Reinforcement Fine-Tuning
where $\alpha$ is the importance weight (set to 1 for simplicity), and the $r_{k}^{\rm prg}$ is calculated through Theorem 4.2. Through Eq.5, L2T optimizes the policy $\pi_{\theta}$ to achieve our goal of boosting effectiveness and efficiency in two parts: (i) The first part ensures the correctness of the final answer through the outcome reward $r(x,z_{0:K})$. (ii) The second part introduces the dense process reward $r_{k}^{\rm prg}$, leveraging information-theoretic internal signals to assess the progress of each episode update. It encourages the model to maximize correctness at each step while avoiding redundant information accumulation. Thus, this optimization enables the model to efficiently utilize the limited token budget, progressively improving reasoning effectiveness, and ultimately achieving high-accuracy reasoning outputs.
@@ 5.1 Experimental Settings
We evaluate on multiple reasoning benchmarks, including AIME24-25, AMC, MATH500 [18], MinervaMATH [27], and Omni-MATH [13] (see Appendices D and F for more benchmarks, e.g., code generation). We use DeepScaleR-1.5B-Preview and DeepSeek-R1-Distill-Qwen-1.5B as base models, which already generate reasoning traces marked with ‘<think>’. We compare L2T against (i) outcome reward-based RL methods (e.g., GRPO [41] for deepseek-model family, more in appendices) and (ii) test-time-compute-focused methods, e.g., length penalty [2] and process-reward approaches such as ReST-MCTS [62] and MRT [36]. Since DeepScaleR-1.5B-Preview has already undergone one round of fine-tuning on 40k math question-answer pairs, we fine-tune it on the 919 AIME questions (from 1989 to 2023); for DeepSeek-R1-Distill-Qwen-1.5B, we fine-tune on a random sample of 4,000 question-answer pairs from NuminaMath [28]. Both fine-tuning and evaluation use a maximum token budget of 16,384. For optimization, we set the learning rate to $1e^{-6}$, weight decay to 0.01, and batch size to 256. To approximate the parameter-compression penalty, we employ a one-layer MLP with a Fisher information-matrix damping coefficient of $1\times 10^{-5}$. The hyperparameters $\alpha$ (Eq.5) and $\beta$ (Eq.3) are set to 0.8 and 0.6, respectively. All experiments are run on the A100 GPU clusters. More details of implementation and hyperparameters are provided in Appendix E.
@@
  +outcome-reward RL (GRPO) 44.5 (+1.7) 39.3 (+2.6) 81.5 (-1.5) 84.9 (-0.3) 24.7 (+0.1) 55.0 (+0.5)
  +length penalty 40.3 (-2.5) 30.3 (-6.4) 77.3 (-5.7) 83.2 (-2.0) 23.0 (-1.6) 50.8 (-3.7)
@@ 5.2 Effectiveness and Efficiency Analysis
We compare L2T with the baselines across all the benchmarks and base models, recording both pass@1 accuracy and the required token budget. To reduce variance from limited samples, we use 20 outputs per question. Table 1 and Figure 2 show that L2T attains SOTA performance, achieving the highest reasoning effectiveness with the smallest token budget. For example, compared to outcome-reward-based methods, L2T delivers over a 3.7% gain in pass@1 and roughly doubles token efficiency; compared to methods focused on test-time compute, it achieves more than a 2% accuracy improvement while reducing the token budget by 20%. Moreover, L2T consistently outperforms baselines on multiple datasets with distributions different from the training data, further demonstrating its effectiveness across diverse tasks. These results confirm the superiority of our approach, which achieves effective reasoning with higher efficiency.
@@ 5.3 Ablation study
The effect of different components.
We evaluate three alternative configurations: (i) replacing the fitting information gain with a task-specific reward model; (ii) removing the parameter-compression penalty; and (iii) substituting the low-rank approximation with random sampling of 30% layers. Notably, the overall contribution of our reward has already been demonstrated in Subsection 5.2. From Figure 5 and Appendix F.3, we observe that both the fitting information gain and the parameter-compression penalty are critical for LLM reasoning; although random sampling is faster than low-rank approximation, it introduces additional error. These findings underscore the advantages of our design.
Parameter sensitivity.
We determine the hyperparameters of L2T by evaluating reasoning performance across benchmarks. Both $\alpha$ and $\beta$ are swept over $[0.3,0.9]$. We first use grid search to screen the parameters with a difference of 0.05, then refine it with 0.01, recording the average outcome. As shown in Figure 5, the optimal setting is $\alpha=0.8$ and $\beta=0.6$, also our choices.